# Modeling Item-Level Dynamic Variability with Residual Diffusion for Bundle Recommendation

### v3

@@ Abstract
Existing solutions for bundle recommendation (BR) have achieved remarkable effectiveness for predicting the user’s preference for prebuilt bundles. However, bundle-item (B-I) affiliation will vary dynamically in real scenarios. For example, a bundle themed as ‘casual outfit’ may add ‘hat’ or remove ‘watch’ due to factors such as seasonal variations, changes in user preferences or inventory adjustments. Our empirical study demonstrates that the performance of mainstream BR models may fluctuate or decline under item-level variability. This paper makes the first attempt to address the above problem and proposes a novel Residual Diffusion for Bundle Recommendation (RDiffBR) as a model-agnostic generative framework which can assist a BR model in adapting this scenario. During the initial training of the BR model, RDiffBR employs a residual diffusion model to process the item-level bundle embeddings which are generated by the BR model to represent bundle theme via a forward-reverse process. In the inference stage, RDiffBR reverses item-level bundle embeddings obtained by the well-trained bundle model under B-I variability scenarios to generate the effective item-level bundle embeddings. In particular, the residual connection in our residual approximator significantly enhances BR models’ ability to generate high-quality item-level bundle embeddings. Experiments on six BR models and four public datasets from different domains show that RDiffBR improves the performance of Recall and NDCG of backbone BR models by up to 23%, while only increased training time about 4%.
@@ Introduction
In recent years, lots of works have focused on bundle recommendation and achieved good results. Including factorization models represented by LIRE (LIRE), EFM (EFM), attention-based models represented by DAM (DAM), Attlist (AttList) and GNN-based models represented by CrossCBR (CrossCBR), BundleGT (BundleGT) and DCBR (DCBR). However, our experiments have found that the performance of existing mainstream bundle recommendation models will fluctuate or even decline under B-I dynamic variability scenarios we have mentioned above. For instance, we found that on Youshu dataset, Recall@20 of BGCN will decrease by up to 19% as items in bundles increases or decreases (see Section 3.2 for details). This is because existing bundle recommendation models lack the ability to effectively model the item-level dynamic variability when generating bundle representations regarding item-level.
@@
Bundle Recommendation. Bundle recommendation aims to recommend a bundle of items related to a specific theme. Early work employed neural network methods for bundle recommendation (EFM; DAM; AttList; LIRE). Recently, graph neural networks have become the mainstream technology for bundle recommendation. We can categorize these graph bundle recommendation models based on these three types of representation learning strategies (Survey_bundle). $\bullet$ Unified learning. The unified representation learning strategy aims to learn representations of users and bundles within a unified user-bundle-item graph (POG; UHBR). $\bullet$ Separate learning. The main idea of this strategy is to first perform representation learning and preference prediction upon two views individually, and then fuse two view-specific predictions to obtain the final prediction (BGCN; IHBR; BundleGT; HBGCN; CMRec). $\bullet$ Cooperative learning. This approach learns user and bundle representations from bundle view and item view and uses techniques such as contrastive learning (MIDGN; CrossCBR; MultiCBR; HIDGN), knowledge distillation (HyperMBR; DGMAE) or diffusion (DCBR) to achieve cross views mutual enhancement. However, most of bundle recommendation lack the ability to adapt to item-level dynamic variability scenario. It is necessary to study our task for real-world application scenarios.
Diffusion Models for Recommendation. The emerging diffusion model in generative models, compared with GANs and VAEs, possess stronger stability and representation. Diffusion model plays three main roles in recommendation systems (survey_diffusion): $\bullet$ Diffusion for data augmentation. This approach generating realistic samples to expand the original training data, such as sequence data (Diff4rec; Plug_diff) and side information (DiffKG; DiffMM). $\bullet$ Diffusion for representation enhancement. Diffusion models converts the input data into embeddings to address some key challenges of recommendation. Lots of studies have been proposed such as denoising the implicit user feedback (DDRM), sequential recommendation (DiffuRec), out-of-distribution issue (CausalDiffRec) and multi-modal learning (MCDRec). $\bullet$ Diffusion as recommendation model. This approach can be classified into three categories: collaborative recommendation (DRM; CFDiff), context-aware recommendation (RecDiff; Diff-POI), and other applications (DCDR). However, a straightforward application of diffusion models on bundle recommendation is insufficient to address the issue of item-level dynamic variability. There are mainly two reasons: the loss targets of diffusion and bundle recommendation are inconsistency and the approximator needs to be designed base on item-level dynamic variability scenario.
@@
Output: A well-trained bundle recommendation estimates the probability $\hat{y}_{ub}\in(0,1)$ of a user $u$ interacting with unseen bundles $b$ based on $Z^{*}=Z\cup\Delta^{+}\setminus\Delta^{-}$ rather than $Z$. Here, $\Delta^{+}=\{(b,i)\mid(b,i)\notin Z\}$ and $\Delta^{-}=\{(b,i)\mid(b,i)\in Z\}$ represent dynamic variability bundle-item affiliation. And when an item is replaced by others, this can be treated as the joint operation of $\Delta^{+}$ and $\Delta^{-}$.
@@
$\tilde{\mu}(e_{t},e_{0},t)$ and $\sigma^{2}(t)I$ are the mean and covariance of $q(e_{t-1}|e_{t},e_{0})$. Following  (DDRM) we can similarity factorize $p_{\theta}(e_{t-1}|e_{t})$:
@@ 13     Take gradient descent step on $\nabla_{\theta}(\mathcal{L})$ to optimize $\theta$;
In order to evaluate the additional computing time of RDiffBR, we compared the one-epoch training time (seconds) of using BR model and BR model+RDiffBR on four datasets (The training rounds are the same for both modes). We recorded five consecutive training epochs and averaged them to obtain the one-epoch training time, as shown in Table 4. It shows that RDiffBR only increased training time about 4%.
@@
In this work, to address the issue of item-level dynamic variability, we proposed RDiffBR, a model-agnostic generative framework that can help a BR model adapt to this scenario. A number of experimental results show that RDiffBR enables backbone BR models to achieve significant performance improvement under dynamic changes of B-I. Future research will consider extending RDiffBR as a plug-and-play module for well-performing pre-trained bundle recommenders and evaluating its applicability to cold-start bundle recommendation.

### v4

@@ Abstract
Existing solutions for bundle recommendation (BR) have achieved remarkable effectiveness for predicting the user’s preference for prebuilt bundles. However, bundle-item (B-I) affiliation will vary dynamically in real scenarios. For example, a bundle themed as ‘casual outfit’ may add ‘hat’ or remove ‘watch’ due to factors such as seasonal variations, changes in user preferences or inventory adjustments. Our empirical study demonstrates that the performance of mainstream BR models may fluctuate or decline under item-level variability. This paper makes the first attempt to address the above problem and proposes Residual Diffusion for Bundle Recommendation (RDiffBR) as a model-agnostic generative framework which can assist a BR model in adapting this scenario. During the initial training of the BR model, RDiffBR employs a residual diffusion model to process the item-level bundle embeddings which are generated by the BR model to represent bundle theme via a forward-reverse process. In the inference stage, RDiffBR reverses item-level bundle embeddings obtained by the well-trained bundle model under B-I variability scenarios to generate the effective item-level bundle embeddings. In particular, the residual connection in our residual approximator significantly enhances BR models’ ability to generate high-quality item-level bundle embeddings. Experiments on six BR models and four public datasets from different domains show that RDiffBR improves the performance of Recall and NDCG of backbone BR models by up to 23%, while only increases training time about 4%.
@@ Introduction
In recent years, lots of works have focused on bundle recommendation and achieved good results. Including factorization models represented by LIRE (LIRE), EFM (EFM), attention-based models represented by DAM (DAM), Attlist (AttList) and GNN-based models represented by CrossCBR (CrossCBR), BundleGT (BundleGT) and DCBR (DCBR). However, our experiments have found that the performance of existing mainstream bundle recommendation models will fluctuate or even decline under B-I dynamic variability scenarios we have mentioned above. For instance, we found that on Youshu dataset, Recall@20 of BGCN will decrease by up to 19% as items in bundles increases or decreases (see Section 3.2 for details). This is because existing bundle recommendation models lack the ability to effectively model the item-level dynamic variability when generating bundle representations of item-level.
@@
Bundle Recommendation. Bundle recommendation aims to recommend a bundle of items related to a specific theme. Early work employed neural network methods for bundle recommendation (EFM; DAM; AttList; LIRE). Recently, graph neural networks have become the mainstream technology for bundle recommendation. We can categorize these graph bundle recommendation models based on these three types of representation learning strategies (Survey_bundle). $\bullet$ Unified learning. The unified representation learning strategy aims to learn representations of users and bundles within a unified user-bundle-item graph (POG; UHBR). $\bullet$ Separate learning. The main idea of this strategy is to first perform representation learning and preference prediction upon two views individually, and then fuse two view-specific predictions to obtain the final prediction (BGCN; IHBR; BundleGT; HBGCN; CMRec). $\bullet$ Cooperative learning. This approach learns user and bundle representations from bundle view and item view and uses techniques such as contrastive learning (MIDGN; CrossCBR; MultiCBR; HIDGN), knowledge distillation (HyperMBR; DGMAE) or diffusion (DCBR) to achieve cross views mutual enhancement. However, most of bundle recommendation lack the ability to adapt to item-level dynamic variability. It is necessary to study our task for real-world application scenarios.
Diffusion Models for Recommendation. The emerging diffusion model in generative models, compared with GANs and VAEs, possess stronger stability and representation. Diffusion model plays three main roles in recommendation systems (survey_diffusion): $\bullet$ Diffusion for data augmentation. This approach generating realistic samples to expand the original training data, such as sequence data (Diff4rec; Plug_diff) and side information (DiffKG; DiffMM). $\bullet$ Diffusion for representation enhancement. Diffusion models converts the input data into embeddings to address some key challenges of recommendation. Lots of studies have been proposed such as denoising the implicit user feedback (DDRM), sequential recommendation (DiffuRec), out-of-distribution issue (CausalDiffRec) and multi-modal learning (MCDRec). $\bullet$ Diffusion as recommendation model. This approach can be classified into three categories: collaborative recommendation (DRM; CFDiff), context-aware recommendation (RecDiff; Diff-POI), and other applications (DCDR). However, a straightforward application of diffusion models on bundle recommendation is insufficient to address the issue of item-level dynamic variability. There are mainly two reasons: the loss targets of diffusion and bundle recommendation are inconsistent and the approximator needs to be designed based on item-level dynamic variability scenario.
@@
Output: A well-trained bundle recommendation estimates the probability $\hat{y}_{ub}\in(0,1)$ of a user $u$ interacting with unseen bundles $b$ based on $Z^{*}=Z\cup\Delta^{+}\setminus\Delta^{-}$ rather than $Z$. Here, $\Delta^{+}=\{(b,i)\mid(b,i)\notin Z\}$ and $\Delta^{-}=\{(b,i)\mid(b,i)\in Z\}$ represent dynamic variability bundle-item affiliation. When an item is replaced by others, this can be treated as the joint operation of $\Delta^{+}$ and $\Delta^{-}$.
@@
$\tilde{\mu}(e_{t},e_{0},t)$ and $\sigma^{2}(t)I$ are the mean and covariance of $q(e_{t-1}|e_{t},e_{0})$. Following  (DDRM) we can similarly factorize $p_{\theta}(e_{t-1}|e_{t})$:
@@ 13     Take gradient descent step on $\nabla_{\theta}(\mathcal{L})$ to optimize $\theta$;
In order to evaluate the additional computing time of RDiffBR, we compare the one-epoch training time (seconds) of using BR model and BR model+RDiffBR on four datasets (The training rounds are the same for both modes). We record five consecutive training epochs and averaged them to obtain the one-epoch training time, as shown in Table 4. It shows that RDiffBR only increased training time about 4%.
@@
In this work, to address the issue of item-level dynamic variability, we propose RDiffBR, a model-agnostic generative framework that can help BR models adapt to this scenario. A number of experiments show RDiffBR enables backbone BR models to achieve significant performance improvement under dynamic changes of B-I. Future research will consider extending RDiffBR as a plug-and-play module for well-performing pre-trained bundle recommenders and evaluating its applicability to cold-start bundle recommendation.